# Proactive Agents for Multi-Turn Text-to-Image Generation Under Uncertainty

### v2

@@ 1 Introduction
At the core of these agents, we build in a graph-based symbolic belief state, termed the belief graph, that allows agents to express uncertainty about entities (e.g., “breakfast”, “table”) that may appear in the image, attributes of entities (e.g., cuisine type) and relations between entities (e.g., “breakfast” being on “table”). When a user provides a prompt, we use an LLM and constrain its generation to adhere to the belief graph data structure, which includes probability estimates for entity presence and potential values for attributes and relations.
For example, for a prompt related to “homemade breakfast on a wooden table” in Figure 1, the agent’s belief graph might indicate a 9% probability that the cuisine is American, while also recognizing other possibilities like Indian or Faroese. The belief graph also anticipates entities not explicitly mentioned but the user might want. Even though the user only specifies “homemade breakfast”, the agent might be 35% confident that a fork will be in the generated images. The expression of uncertainty enables the agent to inspire new ideas and to guide users in editing uncertain items within the graph.
@@
To address the lack of evaluation standards for proactive T2I agents, we develop a scalable and automatic evaluation pipeline with a collection of metrics. The system first constructs a simulated user whose ground truth intent is a prefixed image, and translates the image to a short, underspecified text description. We then let a proactive agent and the simulated user play a multi-turn question answering game. The proactive agent aims to collect as much information as possible to generate images close to the ground truth, while the simulated user provides concise answers to agent questions. We then use text and image similarity metrics to evaluate the generated image against the goal image. We observe a strong correlation between the automatic evaluation results and human evaluation results, highlighting our pipeline’s reliability in assessing T2I agents.
@@ 3 Background
The belief graph in our work is closely related to symbolic world representations. More extensive related works on human-computer interaction, interpretability, prompt expansion are in §2.
World states.  Classical AI represents the world with symbols (McCarthy and Hayes, 1969; Minsky, 1974, 1988; Pasula et al., 2007; Kaelbling and Lozano-Pérez, 2011). For example, in the blocks world (Ginsberg and Smith, 1988; Gupta and Nau, 1992; Alkhazraji et al., 2020), a state can be $is\_block(a)\wedge is\_red(a)\wedge on\_table(a)\wedge is\_block(b)\wedge is\_blue(b)\wedge on(b,a),$ describing that there are a red block and a blue block, referred to as $a$ and $b$, block $a$ is on a table, and block $b$ is on $a$. Such world states include entities (e.g., $a$ and $b$), their attributes (e.g., position $on\_table$, characteristics $is\_block$) and relations (e.g., $on(b,a)$) which are critical for enabling a robot to know and act in the world. The predicates like $on\_table$ are pre-defined and hardcoded into robot systems.
@@
Belief states.  Term “belief state” (Nilsson, 1986; Kaelbling et al., 1998) has been used to describe a distribution over states. E.g., for block $a$, we might have $p(on\_table(a))=0.5$ and $p(\neg on\_table(a))=0.5$, which means the agent is unsure whether the block is on a table. To represent the T2I agent’s belief on which image to generate, we need to consider the distribution over all possible “worlds” in which the picture can be taken. This distribution can be described by the probabilities that an entity appears in the picture, an attribute gets assigned a certain value, etc.
In contrast to classical belief states, our belief graphs incorporate importance scores tailored to generative AI tasks and leverage LLMs to automatically derive predicates from user descriptions. This approach makes the belief graphs highly adaptable to new tasks, eliminating the need for the rigid, hardcoded rules in traditional AI systems.
@@ 4 Belief Graphs
We have 3 types of entities: (a) explicit: entities mentioned in the prompt, (b) implicit: entities not mentioned in the prompt but likely to appear, e.g., fork in fig. 1, and (c) background entities, such as image style, time of day, location, which play important roles in constructing the image.
@@
An entity’s importance score (normalized between 0 and 1) reflects its relevance and impact on the image. For example, in the “breakfast spread out across a wooden table” prompt in fig. 1, we care more about the specific breakfast items than the type of wood the table is made of. Therefore, “breakfast” should have a higher importance score than “table”. It is worth noting that different people have different judgments on the relative importance of elements in an image. In this work, we make a simplifying assumption that the agent doesn’t consider the distribution over people’s perceptions of importance scores.
An entity has multiple attributes, each with a name, a list of value-probability pairs (see “cuisine” in fig. 1) and an importance score. An attribute’s importance score shows its relevance to the image. For example, the cuisine of the breakfast might be more important than its exact size.
@@ 5 Proactive T2I Agent Design
We provide high-level principles and design that guide our agent how to behave and interact with users to generate desired images from text through multi-turn interactions. The goal of the agent is to generate images that match the user’s intended image as closely as possible with minimal back-and-forth, particularly in cases with underspecified prompts and the agent needs to gather information proactively. This requires a decision strategy on information gathering to trade off between the cost of interactions and the quality of generated images. The formal problem definition can be found in §B. Algorithm 1 summarizes how the agent constructs an initial belief graph from user prompt, takes actions (including asking questions and presenting belief graphs), incorporates user feedback during interaction, and updates the belief graph and prompt.
@@
1:  Input: Initial Prompt ($p$)
@@
4:  Parse entity attributes and relations from entities and prompt $p$ (D.6, D.7)
@@
6:  for $turn\leftarrow 1$ to $max\_turn$ do
@@ 5.1 Belief Graph Transitions
The agent’s belief graph gets updated when the agent receives new information through user feedback. This feedback can come from agent-user dialog or through a user direct manipulation of the belief graph on the agent interface (Figure 1). The transition process integrates information from the initial user prompt, the conversation history, interaction and the previous belief. In an agent prototype, we use a simple approach to implement the transition: Generate a comprehensive prompt that summarizes all interactions and information gathered thus far. This merged prompt is then used to re-generate the belief, effectively incorporating the new information into a refreshed representation. The implementation details can be found in §D.4.
@@ 5.2 Asking Questions to Gather Information
The Easy-to-Answer principle focuses on decreasing the cognitive load on users when they are asked to answer questions. A common strategy is for the agent to offer pre-defined answer choices. These choices are typically the ones the agent deems most probable. For example, What color is the cat? (a) Black (b) Brown (c) Orange (d) Other (please specify).
@@ 6 Implemented Agent Prototypes
We propose and experiment with three agent prototypes (Ag1, Ag2, and Ag3). All agents use the same belief graph construction approach in Algorithm 1. They differ in how they address the four question-asking principles (detailed in §5.2), specifically through their decision strategy $\pi$.
@@ 6.1 Ag1: Heuristic Score Agent
Ag1 asks questions about an entity’s existence, attribute values, or relation types when these elements are both highly important and uncertain. It calculates a heuristic score in Equation 1 by multiplying importance scores and entropy of every belief graph element, then selects the highest-scoring question. We then prompt an LLM (§D.12) to craft easy-to-answer, multiple-choice questions like, “What color of the rabbit do you have in mind? a. black, b. white, c. brown, d. unknown, or something else?”
For belief graph transition, Ag1 re-generates the belief graph based on the merged prompt ($p$ in Line 9 of Algorithm 1), and applies the post-processing logic outlined in §D.4 to ensure consistency and prevent redundancy.
Figure 4 shows Ag1 excels at constructing easy-to-answer questions. However, its hardcoded post-processing and heuristic function doesn’t consistently prevent redundancy, largely due to potential LLM errors in parsing beliefs from prompts. For example in Figure 5, Ag1 asks whether “cake” exists, even when it was mentioned in the initial prompt. This can happen if the LLM assigns “cake” a less than 100% existence probability and a high importance score.
@@ 6.2 Ag2: Belief-prompted Agent
Ag2 does not require post-processing logic for the belief transition, since the LLM is expected to use the conversation history to judge redundancy. Our experiments in §7.2 support this hypothesis, showing that Ag2 tends to achieve better performance than Ag1 with fewer turns of questions. However, Ag2 generates fewer easy-to-answer questions than Ag1, as shown in Figure 4.
@@ 6.3 Ag3: Principle-prompted Agent
Ag3 tends to ask more open-ended questions such as “What else is present in the image?” and more complex questions like “Is the cake on a plate or a stand, and what is its color and shape?” Figure 4, Figure 3 and Table 1 show that these types of questions are more difficult for users to answer but gain more information about user intent.
@@ 7.1 Automatic Evaluation
We simulate user-agent conversations using LLM self-play in Algorithm 2, where the conversation begins with a goal image and a detailed ground truth prompt describing it. We use the algorithm similar to Ag2 to simulate the user, answering questions based on the ground truth prompt and the belief graph generated from the ground truth prompt. More details of the simulated user can be found in §D.3. We run each agent-user conversation for 15 turns and compute various metrics at the end of each turn. Figure 2 part b shows the multi-turn set up that we use in our results.
@@ 7.1.1 Setups for agents and baseline
Model Selection. We implement the agent belief parsing and interaction in Algorithm 1 on top of the Gemini 1.5 (Gemini Team Google, 2024) using the default temperature and a 32K context length. For T2I generation, we use Imagen 3 (Baldridge et al., 2024) across all baselines given its recency and prompt-following capabilities. We used both the models served publicly using the Vertex API. [^3] https://cloud.google.com/vertex-ai
@@
The evaluation data used consists of tuples: $(\mathbf{I},p_{0},c,b_{gt})$, where $\mathbf{I}$ represents the target image, $p_{0}$ is an initial prompt describing only the primary elements of the scene, $c$ is a ground truth caption providing a detailed description of $\mathbf{I}$, including spatial layout, background elements, and style, and $b_{gt}$ is the ground truth belief graph constructed via parsing $c$. The initial prompt $p_{0}$ is intentionally less detailed than $c$ to necessitate multi-turn refinement. The evaluation framework we propose directly assesses the agent’s ability to guide the user towards the target image $\mathbf{I}$ starting from a $p_{0}$.
Existing image-caption datasets primarily focus on simple scenes (Deng et al., 2009; Krizhevsky et al., 2009; Deng, 2012) or focus on very specific categories (Liu et al., 2016; Liao et al., 2022). With the aim for complex realistic images for testing the robustness of the agents, we evaluate over the Coco-Captions dataset validation split (Chen et al., 2015) and ImageInWords (Garg et al., 2024) dataset. ImageInWords dataset contains a diverse set of realstic and cartoon images and has human annotators create dense detailed captions that describe attributes and relationships between objects in the image. In ImageInWords evaluations we use the long human annotation as the ground truth caption and auto-summarize the long caption with Gemini 1.5 Pro to create starting prompt $p_{0}$. For each image in Coco-Captions, the dataset provides multiple human generated captions which are short and describe the basic elements and the interactions in the image. We select the shortest of captions per image as the starting prompt $p_{0}$. We obtain the ground truth caption by expanding $p_{0}$ via prompting Gemini 1.5 Pro with ground truth image and asking it to add more details of the attributes of the entities in the image as well as the style and image composition.
COCO-Captions and ImageInWords datasets lack artistic and non-photorealistic imagery often desired by designers and artists. To better evaluate our target for flexible use cases such as by artists, we introduce and release DesignBench, a novel dataset comprising 30 scenes specifically designed for this purpose. Each scene follows the $(\mathbf{I},p_{0},c,b_{gt})$ format described earlier. DesignBench includes a mix of human generated cartoon graphics, photorealistic yet improbable scenes, and artistic photographic images. $p_{0}$ and $c$ in DesignBench are generated via prompting Gemini 1.5 Pro with ground truth image.
@@ 7.1.3 Metrics
Text-Image Similarity: We compare the ground truth prompt with the generated image (T2I) using VQAScore (Lin et al., 2024). We use the author released implementation of the metric and use Gemini 1.5 Pro as the underlying multi-modal large language model.
@@ 7.2 Automatic Metrics and Qualitative Results
Results in Table 1 show the $\mathbf{I}$, $c$ and $b_{gt}$ against each agent’s final generated image, merged prompt and belief graph. All columns except ‘human rating’ show the mean and standard deviation of the similarity metric at the final turn. The blue row shows the baseline T2I method which generates an image directly from the starting prompt $p_{0}$.
Multi-Turn agents show clear advantage: The results in Table 1 show that significant gains in performance come from using proactive multi-turn agents. The blue row shows the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt $p_{0}$. We see that all of the multi-turn agents far exceed the baseline T2I model on both datasets and all metrics. Ag3 (the LLM agent that does not explicitly utilize the belief graph to generate questions) shows superior performance across all metrics. This confirms that the current T2I agents often produce less desirable images given ambiguity in prompts. In Figure 2 we see real outputs of the multi-turn set up with the Ag3 agent. Additionally, we see that the multi-turn agents generally improve in most metrics as they increase the number of interactions. The plots in Figure 3 show the T2T, I2I, T2I and NLL metrics, averaged across all images in the ImageInWords dataset, per turn for 15 turns. Interestingly we see the T2T and the T2I VQA similarity metric seems to plateau or decrease after about 10 interactions, while the I2I scores continue to increase. The NLL metric shows large performance gains of the Ag3 agent in comparison to all other methods. The plots in Figure 11 show the T2T DSG metrics.
LLMs play a significant role: The best performers (Ag2 and Ag3) both query the LLM to generate a question based on contextual information such as the belief graph and conversation history. They query the LLM to construct a concise and clear question but don’t impose further constraints on the question construction. Ag1 provides a programmatic template for how the LLM should construct the question based on its belief graph and does not provide any conversation history information. Examples of dialogs and the generated questions produced by the three agents can be found in the Appendix in Figure 5. This figure demonstrates that the templated question creation sometimes leads to questions that gather minimal information in return. This can be an intrinsic limitation of hard coded question selection strategy but also can be an issue of the heuristic scores we defined for question selection in Ag1. In contrast, Ag2 and Ag3 generate questions that are more open-ended thus allowing the user to provide more nuanced details which in consequence enhance the agent’s image knowledge.
Figure 3: ImageInWords results, including (a) T2T, (b) I2I, (c) T2I, (d) NLL scores. Agents trend to increase performance up to 10 turns.

### v3

@@ 1 Introduction
At the core of these agents, we build in a graph-based symbolic belief state, termed the belief graph, that allows agents to express uncertainty about entities (e.g., “breakfast”, “table”) that may appear in the image, attributes of entities (e.g., cuisine type) and relations between entities (e.g., “breakfast” being on “table”). Given a prompt, we use an LLM and constrain its generation to adhere to the belief graph data structure, which includes probability estimates for entity presence and potential values for attributes and relations.
For example, for a prompt related to “homemade breakfast on a wooden table” in Figure˜1, the agent’s belief graph might indicate a 9% probability that the cuisine is American, while simultaneously recognizing other possibilities like Indian or Faroese. The belief graph also anticipates entities not explicitly mentioned but the user might want. Even though the user only specifies “homemade breakfast”, the agent might be 35% confident that a fork will be in the generated images. The expression of uncertainty enables the agent to inspire new ideas and to guide users in editing uncertain items within the graph.
@@
To address the lack of evaluation standards for proactive T2I agents, we develop a scalable and automatic evaluation pipeline with a collection of metrics. The system first constructs a simulated user whose ground truth intent is a prefixed image, and translates the image to a short, underspecified text description. We then let a proactive agent and the simulated user play a multi-turn question answering game. The proactive agent aims to collect as much information as possible to generate images close to the ground truth, while the simulated user provides concise answers to agent questions. We then use text and image similarity metrics to evaluate the generated image against the goal image.
@@ 3 Background
The belief graph in our work is closely related to symbolic world representations. See §2 for other related works on human-computer interaction, interpretability, etc.
World states.  Classical AI represents the world with symbols (McCarthy and Hayes, 1969; Minsky, 1974, 1988; Pasula et al., 2007; Kaelbling and Lozano-Pérez, 2011). For example, in the blocks world (Ginsberg and Smith, 1988; Gupta and Nau, 1992; Alkhazraji et al., 2020), a state can be $is\_block(a)\wedge is\_red(a)\wedge on\_table(a)\wedge is\_block(b)\wedge is\_blue(b)\wedge on(b,a),$ describing that there are a red block and a blue block, referred to as $a$ and $b$, block $a$ is on a table, and block $b$ is on $a$. Such world states include entities (e.g., $a$ and $b$), their attributes (e.g., position on_tableon\_table, characteristics is_blockis\_block) and relations (e.g., $on(b,a)$) which are critical for enabling a robot to know and act in the world. The predicates like on_tableon\_table are pre-defined and hardcoded into robot systems.
@@
Belief states.  Term “belief state” (Nilsson, 1986; Kaelbling et al., 1998) has been used to describe a distribution over states. E.g., for block $a$, we might have p(on_table(a))=0.5p(on\_table(a))=0.5 and $p(\neg on\_table(a))=0.5$, which means the agent is unsure whether the block is on a table. To represent the T2I agent’s belief on which image to generate, we need to consider the distribution over all possible “worlds” in which the picture can be taken. This distribution can be described by the probabilities that an entity appears in the picture, an attribute gets assigned a certain value, etc.
In contrast to classical belief states, our belief graphs incorporate importance scores tailored to generative AI tasks and leverage LLMs to automatically derive predicates from user prompts. This approach makes the belief graphs highly adaptable to new tasks, eliminating the need for the rigid, hardcoded rules in traditional AI systems.
@@ 4 Belief Graphs
We have 3 types of entities: (a) explicit: entities mentioned in the prompt, (b) implicit: entities not mentioned in the prompt but likely to appear, e.g., “fork” in fig.˜1, and (c) background entities, such as image style, time of day, location, which play important roles in constructing the image.
@@
An entity’s importance score (normalized between 0 and 1) reflects its relevance and impact on the image. For example, in the “breakfast spread out across a wooden table” prompt in fig.˜1, we care more about the specific breakfast items than the type of wood the table is made of. Therefore, “breakfast” should have a higher importance score than “table”. It is worth noting that different people have different judgments on the relative importance of elements in an image. In this work, we make a simplifying assumption that the agent doesn’t consider the distribution over people’s perceptions of importance scores.
An entity has multiple attributes, each with a name, a list of value-probability pairs (e.g., see the cuisine attribute of “breakfast” in fig.˜1) and an importance score. An attribute’s importance score shows its relevance to the image. For example, the cuisine of the breakfast might be more important than its exact size.
@@ 5 Proactive T2I Agent Design
We provide high-level principles and design that guide our agent how to behave and interact with users to generate desired images from text through multi-turn interactions. The goal of the agent is to generate images that match the user’s intended image as closely as possible with minimal back-and-forth, particularly in cases with underspecified prompts and the agent needs to gather information proactively. This requires a decision strategy on information gathering to trade off between the cost of interactions and the quality of generated images. The formal problem definition can be found in §B. Algorithm 1 summarizes how the agent constructs an initial belief graph from a user prompt, takes actions (including asking questions and presenting belief graphs), incorporates user feedback during interaction, and updates the belief graph and prompt.
@@
1:Input: Initial user prompt ($p$)
@@
4: Parse attributes and relations from entities and $p$ (D.6, D.7)
@@
6:for $turn\leftarrow 1$ to max_turnmax\_turn do
@@ 5.1 Belief Graph Transitions
The agent’s belief graph gets updated when the agent receives new information through user feedback. This feedback can come from agent-user dialog or through a user’s direct manipulation of the belief graph on the agent interface (Figure˜1). The transition process integrates information from the initial user prompt, the conversation history, interaction and the previous belief. In an agent prototype, we use a simple approach to implement the transition: Generate a comprehensive prompt that summarizes all interactions and information gathered thus far. This merged prompt is then used to re-generate the belief, effectively incorporating the new information into a refreshed representation. The implementation details can be found in §D.4.
@@ 5.2 Asking Questions to Gather Information
The Easy-to-Answer principle aims to reduce the mental effort required from users when responding to questions. One strategy is for the agent to offer pre-defined answer choices. These choices are typically the ones the agent deems most probable. For example, What color is the cat? (a) Black (b) Brown (c) Orange (d) Other (please specify).
@@ 6 Implemented Agent Prototypes
We propose and experiment with three agent prototypes (Ag1, Ag2, and Ag3). All agents use the same belief graph construction approach in Algorithm˜1. They differ in how they address the four question-asking principles (detailed in §5.2), specifically through their decision strategy $\pi$.
@@ 6.1 Ag1: Heuristic Score Agent
Ag1 asks questions about an entity’s existence, attribute values, or relation types when these elements are both highly important and uncertain. It calculates a heuristic score in Equation˜1 by multiplying importance scores and entropy of every belief graph element, then selects the highest-scoring question. We then prompt an LLM (§D.12) to craft easy-to-answer, multiple-choice questions like, “What color of the rabbit do you have in mind? a. black, b. white, c. brown, d. unknown, or something else?”
For belief graph transition, Ag1 re-generates the belief graph based on the merged prompt ($p$ in Line 9 of Algorithm˜1), and applies the post-processing logic outlined in §D.4 to ensure consistency and prevent redundancy.
Figure˜4 shows Ag1 excels at constructing easy-to-answer questions. However, its hardcoded post-processing and heuristic function doesn’t consistently prevent redundancy, largely due to potential LLM errors in parsing beliefs from prompts. For example in Figure˜5, Ag1 asks whether “cake” exists, even when it was mentioned in the initial prompt. This can happen if the LLM assigns “cake” a less than 100% existence probability and a high importance score.
@@ 6.2 Ag2: Belief-prompted Agent
Ag2 does not require the post-processing logic (§D.4) for the belief transition, since the LLM is expected to use the conversation history to judge redundancy. Our experiments in §7.2 support this hypothesis, showing that Ag2 tends to achieve better performance than Ag1 with fewer turns of questions. However, Ag2 generates fewer easy-to-answer questions than Ag1, as shown in Figure˜4.
@@ 6.3 Ag3: Principle-prompted Agent
Ag3 tends to ask more open-ended questions such as “What else is present in the image?” and more complex questions like “Is the cake on a plate or a stand, and what is its color and shape?” Figure˜4, Figure˜3 and Table˜1 show that these types of questions are more difficult for users to answer but gain more information about user intent.
@@ 7.1 Automatic Evaluation
We simulate user-agent conversations using LLM self-play in Algorithm˜2, where the conversation begins with a goal image and a detailed ground truth prompt describing it. We use the algorithm similar to Ag2 to simulate the user, answering questions based on the ground truth prompt and the belief graph generated from the ground truth prompt. More details of the simulated user can be found in §D.3. We run each agent-user conversation for 15 turns and compute various metrics at the end of each turn. Figure˜2 part b shows the multi-turn set up that we use in our results.
@@ 7.1.1 Setups for agents and baseline
Model Selection. We implement the agent belief parsing and interaction in Algorithm˜1 on top of the Gemini 1.5 (Gemini Team Google, 2024) using the default temperature and a 32K context length. For T2I generation, we use Imagen 3 (Baldridge et al., 2024) across all baselines given its recency and prompt-following capabilities. We used both the models served publicly using the Vertex API. [^3] https://cloud.google.com/vertex-ai
@@
The evaluation data consists of tuples: $(\mathbf{I},p_{0},c,b_{gt})$, where $\mathbf{I}$ represents the target image, $p_{0}$ is an initial prompt describing only the primary elements of the scene, $c$ is a ground truth caption providing a detailed description of $\mathbf{I}$, including spatial layout, background elements, and style, and $b_{gt}$ is the ground truth belief graph constructed via parsing $c$. The initial prompt $p_{0}$ is intentionally less detailed than $c$ to necessitate multi-turn refinement. The evaluation framework we propose directly assesses the agent’s ability to guide the user towards the target image $\mathbf{I}$ starting from a $p_{0}$.
Existing image-caption datasets primarily focus on simple scenes (Deng et al., 2009; Krizhevsky et al., 2009; LeCun et al., 2010) or focus on very specific categories (Liu et al., 2016; Liao et al., 2022). With the aim for complex realistic images for testing the robustness of the agents, we evaluate over the Coco-Captions dataset validation split (Chen et al., 2015) and ImageInWords (Garg et al., 2024) dataset. ImageInWords dataset contains a diverse set of realistic and cartoon images and has human annotators create dense detailed captions that describe attributes and relationships between objects in the image. In ImageInWords evaluations we use the long human annotation as the ground truth caption and auto-summarize the long caption with Gemini 1.5 Pro to create starting prompt $p_{0}$. For each image in Coco-Captions, the dataset provides multiple human generated captions which are short and describe the basic elements and the interactions in the image. We select the shortest of captions per image as the starting prompt $p_{0}$. We obtain the ground truth caption by expanding $p_{0}$ via prompting Gemini 1.5 Pro with ground truth image and asking it to add more details of the attributes of the entities in the image as well as the style and image composition.
COCO-Captions and ImageInWords datasets lack artistic and non-photorealistic imagery often desired by designers and artists. To better evaluate our target for flexible use cases such as by artists, we introduce and release DesignBench [^4] https://huggingface.co/datasets/meerahahn/DesignBench, a novel dataset comprising 30 scenes specifically designed for this purpose. Each scene follows the $(\mathbf{I},p_{0},c,b_{gt})$ format described earlier. DesignBench includes a mix of human generated cartoon graphics, photorealistic yet improbable scenes, and artistic photographic images. Initial prompt $p_{0}$ and ground truth caption $c$ in DesignBench are generated via prompting Gemini 1.5 Pro with ground truth image.
@@ 7.1.3 Metrics
Text-Image Similarity (T2I): We compare the ground truth prompt with the generated image using VQAScore (Lin et al., 2024). We use the author released implementation of the metric and use Gemini 1.5 Pro as the underlying multi-modal large language model.
@@ 7.2 Automatic Metrics and Qualitative Results
Results in Table˜1 show the $\mathbf{I}$, $c$ and $b_{gt}$ against each agent’s final generated image, merged prompt and belief graph. All columns except ‘human rating’ show the mean and standard deviation of the similarity metric at the final turn. The blue row shows the baseline T2I method which generates an image directly from the starting prompt $p_{0}$.
Multi-Turn agents show clear advantage: The results in Table˜1 show that significant gains in performance come from using proactive multi-turn agents. The blue row shows the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt $p_{0}$. We see that all of the multi-turn agents far exceed the baseline T2I model on both datasets and all metrics. Ag3 (the LLM agent that does not explicitly utilize the belief graph to generate questions) shows superior performance across all metrics. This confirms that the current T2I agents often produce less desirable images given ambiguity in prompts. In Figure˜2 we see real outputs of the multi-turn set up with the Ag3 agent. Additionally, we see that the multi-turn agents generally improve in most metrics as they increase the number of interactions. Figure˜3 shows the T2T, I2I, T2I and NLL metrics, averaged across all images in the ImageInWords dataset, per turn for 15 turns. Interestingly we see the T2T and the T2I VQA similarity scores seem to plateau or decrease after about 10 interactions, while the I2I scores continue to increase. The NLL metric shows large performance gains of the Ag3 agent in comparison to all other methods. Figure˜11 shows the T2T DSG metrics.
LLMs play a significant role: The best performers (Ag2 and Ag3) both query the LLM to generate a question based on contextual information such as the belief graph and conversation history. They query the LLM to construct a concise and clear question but don’t impose further constraints on the question construction. Ag1 provides a programmatic template for how the LLM should construct the question based on its belief graph and does not provide any conversation history information. Examples of dialogs and the generated questions produced by the three agents can be found in the Appendix in Figure˜5. This figure demonstrates that the templated question creation sometimes leads to questions that gather minimal information in return. This can be an intrinsic limitation of hard coded question selection strategy but also can be an issue of the heuristic scores we defined for question selection in Ag1. In contrast, Ag2 and Ag3 generate questions that are more open-ended thus allowing the user to provide more nuanced details which in consequence enhance the agent’s image knowledge.
Figure 3: ImageInWords results, including (a) T2T, (b) I2I, (c) T2I, (d) NLL scores. Agents outperform standard setups of T2I and trend to increase performance up to 10 turns.
@@
7.3 Human Studies on Generated Images and Dialogues
We conducted human studies to provide more insights that complement the quantitative results from our automatic evaluations. See detailed design in Figure˜19, Figure˜20 and Figure˜21. Information on human user recruitment and consent can be found in §H.
Table˜1 shows the human rater evaluations on the generated images in comparison to the ground truth image. Participants are asked to rank the images produced by the three proposed multi-turn agents and a single-turn T2I model in terms of content and aesthetics/style against a ground truth image for which the original prompt was derived and the answers to the agent’s questions were derived. The human ratings correspond to the I2I (DINO) metric. Approximately 550 image-dialog pairs per agent are rated using 3 human raters. The generated unlabeled images were presented in a random order and the human rater was tasked with ranking the images from best to worst in terms of content correctness. The results are in Table 1, under the column Human Rating, showing that agentic systems are selected as the best generated image over the single-turn T2I in 80%+ of cases for both content and aesthetics/style.
Figure˜4 shows the human study results on the generated agent-simulated user dialogues. (1) Per question raters are asked to mark issues each agent question contains that could pose a disturbance to the user, such as ‘question is too long and complex’. Approximately 8k questions per agent are rated. The results in Figure˜4 (left), show agentic systems have issues with their questions in 14% or less cases. Results show Ag2 and Ag3 suffer from too long questions while Ag1 suffers from questions that do not gain new information. (2) Raters are asked to evaluate the correspondence of each image to the agent-user dialog and original prompt. Approximately 1.5k image-dialog pairs are rated using 3 human raters. Results in Figure˜4 (right) show that for all of our agents, more than 96% of the 1.5k image-dialog pairs are rated as very close or fairly close with some differences. This high rating shows the viability of the T2I model employed by the agents, as well as the agents’ ability to combine the dialogue into a coherent prompt to feed the T2I model.